OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]? 312 70 168 −1

OpenStudy (ajprincess):

First find g(5). Put 5 in place of x in g(x). Can u plz tell me vat u get @Solmyr?

OpenStudy (anonymous):

g(5) = (5)2 -1 g(5) = 25-1 g(5) = 24

OpenStudy (ajprincess):

ya perfect:)

OpenStudy (ajprincess):

Nw to find f(g(5)) put g(5) in place of x in f(x). Can u plz tell me vat u get?

OpenStudy (anonymous):

f[g(5)] = 3(5) - 2 f[g(5)] = 15 - 2 f[g(5)] = 13

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

hmm no. u hav to put 24 for x nt 5.

OpenStudy (anonymous):

so f[g(24)] = 13

OpenStudy (ajprincess):

oops no f(g(5))=3(24)-2

OpenStudy (anonymous):

oh f[g(5)] = 3(24) - 2 f[g(5)] = 72 - 2 f[g(5)] = 70

OpenStudy (ajprincess):

ya perfect:)

OpenStudy (anonymous):

so is the answer 70?

OpenStudy (anonymous):

B

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

ok thank you

OpenStudy (ajprincess):

yw:)